OpenStudy (anonymous):

Explain how applying the Commutative Property of Multiplication to twenty-five times twenty-two times 4 makes it easier to simplify. Show the solution in your explanation.

OpenStudy (mertsj):

Do you know what 25 times 4 is?

OpenStudy (mertsj):

I guess not.

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

\[25\times 22\times 4\]?

OpenStudy (anonymous):

are you asking what they are?

OpenStudy (anonymous):

easier if you rewrite it as \[25\times 4\times 22\] because \(25\times 4=100\) and it is simple to multiply by \(100\)

OpenStudy (anonymous):

rewrite what mate?

OpenStudy (anonymous):

\[25\times 22\times 4\] as \[25\times 4\times 22\] via the commutative property of multiplication

OpenStudy (anonymous):

ohh okay

OpenStudy (the_fizicx99):

Commutative property of multiplication means no matter what way you put it, its still going to give you the same result. i.e 22 * 25 * 4 will be the same as 4 * 25 * 22 So by multiplying 25 * 4 = 100 * 22 will no change your result in anyway

OpenStudy (anonymous):

really?

OpenStudy (texaschic101):

commutative property basically states that you can move your numbers around and still get the same answer. 25 x 22 x 4....It would be easier to write it like this...25 x 4 x 22 because 25 x 4 = 100 and 100 x 22 = 2200.

OpenStudy (the_fizicx99):

Yes! @stolenaxis43

OpenStudy (the_fizicx99):

This only apply's to adding and multiplying. 2 + 3 = 5 3 + 2 = 5 4*5 = 20 5*4 = 20

OpenStudy (anonymous):

okay now i get it thank you um... lad or lass?

OpenStudy (the_fizicx99):

:)